# SelfElicit: Your Language Model Secretly Knows Where is the Relevant Evidence

### v1

@@ Abstract
Providing Language Models (LMs) with relevant evidence in the context (either via retrieval or user-provided) can significantly improve their ability to provide factually correct grounded responses. However, recent studies have found that LMs often struggle to fully comprehend and utilize key evidence from the context, especially when it contains noise and irrelevant information—an issue common in real-world scenarios. To address this, we propose SelfElicit, an inference-time approach that helps LMs focus on key contextual evidence through self-guided explicit highlighting. By leveraging the inherent evidence-finding capabilities of LMs using the attention scores of deeper layers, our method automatically identifies and emphasizes key evidence within the input context, facilitating more accurate and factually grounded responses without additional training or iterative prompting. We demonstrate that SelfElicit brings consistent and significant improvement on multiple evidence-based QA tasks for various LM families while maintaining computational efficiency. Our code and documentation are available at https://github.com/ZhiningLiu1998/SelfElicit.
SelfElicit: Your Language Model Secretly Knows Where is the Relevant Evidence
Zhining Liu1, Rana Ali Amjad2, Ravinarayana Adkathimar2, Tianxin Wei1, Hanghang Tong1 1University of Illinois Urbana-Champaign, 2Amazon Science 1{liu326, twei10, htong}@illinois.edu 2{raamjad,adkathi,ani}@amazon.com
@@ 1 Introduction
Language models have demonstrated remarkable capabilities OpenAI (2022, 2023); Kaddour et al. (2023), yet hallucinations and factual inaccuracies remain a significant limitation Ji et al. (2023); Wang et al. (2023). Furnishing the context with relevant evidence is a popular approach to enhance LM’s ability to generate factually correct grounded responses Ram et al. (2023); Jiang et al. (2023b).
@@
Recently, improved prompting Zhou et al. (2023) or decoding Shi et al. (2024) methods have been proposed to enhance the focus on context. However, they treat the whole context as a single entity, overlooking the fact that not all information provided in the context is important. In this work, we investigate how to leverage contextual information at a finer granularity to help LMs focus more effectively on the key evidence within the context. Our contributions in this work are as follows:
@@
  • Leveraging this inherent ability, we propose an inference time context augmentation approach, SelfElicit, that highlights the important evidence in the context and improves LM’s ability to provide factually correct grounded responses. SelfElicit is efficient for inference, training-free, and robust to noise and hyper-parameter choice.
  • In a comprehensive study, we show the significant performance improvement that SelfElicit brings across model families and benchmarks. We also provide detailed studies for hyper-parameters in SelfElicit and additional analysis on its ability to elicit relevant evidence robustly from noisy context.
Fig. 1 illustrates the SelfElicit workflow. By identifying and highlighting key evidence within the context, SelfElicit guides the LM towards important information that it might otherwise overlook, leading to more factual responses.
@@ 3.2 Contextual Evidence Highlighting
SelfElicit is summarized in Algorithm 1. Note that SelfElicit is computationally efficient as it only generates one additional token (i.e., line 1 in Alg. 1) for identifying ${\mathcal{S}}_{\text{SE}}$ during inference.
@@ 4 Experiments
  • RQ1: How does SelfElicit perform in terms of improving answer quality and factuality?
@@
Llama-3.1 8B Base 58.9 57.7 64.3 56.2 72.8 66.1 59.7 61.6 4.75 4.38 224.1
Cot 60.4 58.6 64.9 55.8 74.4 67.4 59.6 62.1 4.00 3.75 224.8
FullElicit 60.7 59.0 65.9 56.6 72.8 66.3 61.1 62.5 3.25 3.12 226.3
PromptElicit 66.3 66.3 62.8 56.7 76.0 69.6 61.8 65.6 2.00 2.75 1672.0
@@
70B Base 71.8 74.2 66.7 57.4 78.0 71.2 59.3 63.2 3.00 3.25 1389.8
Cot 72.4 74.0 67.2 56.4 77.0 69.9 60.3 63.1 4.50 3.12 1394.2
FullElicit 71.3 73.8 66.2 56.8 77.5 70.7 58.2 61.8 4.50 4.25 1408.1
PromptElicit 73.4 76.2 63.1 58.2 77.0 72.3 64.0 68.5 2.00 3.38 8124.0
@@
PromptElicit 71.0 64.1 57.7 42.2 81.9 62.3 65.6 42.8 1.25 4.00 1877.8
SelfElicit 74.3 61.6 60.6 41.7 83.6 61.3 66.4 43.4 1.75 2.00 431.3
12B Base 59.2 65.9 51.9 51.6 72.9 68.7 54.8 61.6 4.75 5.00 281.9
@@
FullElicit 59.8 65.8 53.1 51.9 73.7 68.8 55.5 62.6 4.00 2.88 283.9
PromptElicit 62.8 73.1 52.2 56.6 79.9 77.6 55.5 65.7 1.75 2.38 1455.0
SelfElicit 63.6 72.9 54.9 58.6 82.6 79.9 55.3 66.0 1.25 1.50 339.1
@@
Cot 70.7 37.9 59.2 31.9 78.6 41.6 63.8 32.3 5.00 2.00 421.5
FullElicit 65.5 65.7 57.5 48.2 77.1 67.3 64.4 60.6 2.75 3.50 249.6
PromptElicit 64.7 67.7 54.9 46.3 75.4 66.8 64.5 65.0 2.25 4.25 1165.1
SelfElicit 69.1 71.4 59.6 50.8 78.1 67.8 65.0 64.7 1.25 1.50 289.4
32B Base 71.8 68.4 60.0 44.7 79.0 69.3 62.7 59.3 3.25 3.12 928.2
Cot 71.3 67.1 60.0 43.5 79.5 66.8 59.5 55.3 5.00 3.62 998.6
FullElicit 71.3 68.2 61.6 45.7 78.5 68.8 63.2 58.1 3.75 3.25 936.3
PromptElicit 71.3 74.5 59.0 51.5 78.0 69.9 64.8 68.1 2.00 4.00 5109.8
@@ 4.1 Experimental Setup
  • Cot Wei et al. (2022): Chain-of-thought prompting encourages the model to reason through intermediate steps before reaching a final answer. While CoT promotes step-by-step reasoning based on the evidence in the context, it does not explicitly highlight important information within the context. Therefore, we use it as a natural baseline for validating the advantage of explicit evidence elicitation in SelfElicit. We implement it by adding the Cot prompt "Think step by step to provide the answer." at the end of the instruction.
  • FullElicit: A naive approach that highlights the entire context as important. Comparing with it demonstrates the necessity of fine-grained, sentence-level evidence elicitation.
@@ 4.2 Main Results
  • Cot also does not provide a meaningful and consistent gain which highlights the fact that asking the model to reason carefully does not improve its ability to leverage the relevant evidence while generating the response. Cot results also show variation across LM families where we don’t observe meaningful gains for Llama and Mistral models. For Qwen models Cot generates longer answers with intermediate steps. While this results in a slight increase in EM scores, it also leads to a large drop (up to 27.9) in F1 scores due to the inclusion of redundant information, as well as longer inference times (e.g., 171.9% more inference time for Qwen2.5-7B)
@@ 4.3 Additional Analysis
However, similar to the baseline, the performance of SelfElicit for "distractor" variant worsens meaningfully when compared to SelfElicit performance for the gold context setup, also shown in Figure 3(a). To investigate this we do a deep dive into the evidence elicitation capability of SelfElicit in the presence of substantial noise. Figure 3(b) contrasts the elicit ratio (i.e., the proportion of elicited evidence within the original context) under gold and distractor settings. With gold context, SelfElicit tends to select larger proportion as evidence since most contextual information provided is relevant. By contrast, in the distractor setting where most of the context is irrelevant, the proportion of evidence selected decreases substantially (typically <10%) for same $\alpha=0.5$. Furthermore the proportion of elicited evidence coming from the additional distractor information is only TKeven though the distractor information constitutes a major portion of the context. Therefore, even though SelfElicit highlights the relevant evidence without being distracted by noise, there is still potential to improve how the LM can utilize this highlighted evidence to maintain similar performance as in the case of gold context, a topic for future research.
@@ 5 Related Works
Furnishing LMs with relevant context is an effective way of providing up-to-date external and/or private knowledge Ji et al. (2023); Asai et al. (2023) to help mitigate hallucination and improve response accuracy Huang et al. (2023). Retrieval-augmented Generation (RAG) is a widely adopted paradigm for this purpose Gao et al. (2023); Fan et al. (2024). Despite its popularity, recent studies have pointed out that context retrieved from external sources often contains noise and irrelevant information, leading to confusion for LM Cuconasu et al. (2024); Wu et al. (2024). Motivated by this, we explore how to leverage contextual information more effectively at a finer granularity by highlighting critical information within the context. To the best of our knowledge, we are the first to investigate automated contextual evidence highlighting based on LM internal representations.
@@
Recent studies have explored ways to understand QA factuality by analyzing the internal representations to identify important attention layers Yuksekgonul et al. (2024); Chen et al. (2024b) or heads Halawi et al. (2024) that are crucial for generation correctness or hallucination. However, they primarily focus on how LMs utilize their parametric knowledge in controlled generation Yuksekgonul et al. (2024); Halawi et al. (2024) or closed-book QA Chen et al. (2024b), with little discussion on the role of internal representations when utilizing external non-parametric knowledge in context-based QA. Our work focuses on investigating LM’s internal attention patterns for the context and on how to leverage these patterns to identify critical evidence in the context.
@@ 6 Conclusion
We evaluated the effectiveness of SelfElicit across several open-source LMs. However, we can not assess SelfElicit on proprietary LMs due to the need to access attention scores. On the computational efficiency front, although SelfElicit generates only one additional token in the first pass and already demonstrates good computational efficiency in practice, there is potential for further acceleration by avoiding re-encoding prompt. This redundancy arises because the input sequence is encoded twice during the first and second passes, despite the only difference being a few text markers inserted into the context passage for evidence highlighting. Exploring methods for caching and reusing shared input content encoding process, thereby avoiding the repeated encoding of the input content is a promising future direction for further reducing computational overhead.
@@ Appendix A Reproducibility Details
We use PyTorch Paszke et al. (2019) and HuggingFace Transformers Wolf et al. (2020) to implement all involved language models (LMs). Specifically, the LMs used in the paper are the instruct fine-tuned version of Llama3.1 8B [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct and 70B [^4] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, Mistral 7B (instruct-v0.3) [^5] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 and 12B (Nemo) [^6] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, as well as Qwen2.5 7B [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and 32B [^8] https://huggingface.co/Qwen/Qwen2.5-32B-Instruct. We use the official apply_chat_template function to get the input tokens from text input ($\tau({\mathbf{c}},{\mathbf{q}})$) and let the model generate the answer based on it. We use greedy decoding with a temperature of 0 to get deterministic answers. For all the datasets used, we test the LM on the official public validation split, with 7,405/4,212/7,785/12,836 context-question pairs for HotopotQA/NewsQA/TQA/NQ.

### v2

@@ Abstract
Providing Language Models (LMs) with relevant evidence in the context (either via retrieval or user-provided) can significantly improve their ability to provide better-grounded responses. However, recent studies have found that LMs often struggle to fully comprehend and utilize key evidence from the context, especially when it contains noise and irrelevant information—an issue common in real-world scenarios. To address this, we propose SelfElicit, an inference-time approach that helps LMs focus on key contextual evidence through self-guided explicit highlighting. By leveraging the inherent evidence-finding capabilities of LMs using the attention scores of deeper layers, our method automatically identifies and emphasizes key evidence within the input context, facilitating more accurate and grounded responses without additional training or iterative prompting. We demonstrate that SelfElicit brings consistent and significant improvement on multiple evidence-based QA tasks for various LM families while maintaining computational efficiency. Our code and documentation are available at https://github.com/ZhiningLiu1998/SelfElicit.
SelfElicit: Your Language Model Secretly Knows
Where is the Relevant Evidence
Zhining Liu1, Rana Ali Amjad2, Ravinarayana Adkathimar2, Tianxin Wei1, Hanghang Tong1 1University of Illinois Urbana-Champaign, 2Amazon Science 1{liu326, twei10, htong}@illinois.edu 2{raamjad,adkathi}@amazon.com
@@ 1 Introduction
Language models have demonstrated remarkable capabilities OpenAI (2022, 2023); Kaddour et al. (2023), yet hallucinations and factual inaccuracies remain a significant limitation Ji et al. (2023); Wang et al. (2023a). Furnishing the input context with relevant evidence passage is a popular approach to enhance LM’s ability to generate correct and well-grounded responses Ram et al. (2023); Jiang et al. (2023b); Wei et al. (2025b).
@@
Recently, improved prompting Zhou et al. (2023) or decoding Shi et al. (2024) methods have been proposed to enhance the focus on context. However, they treat the whole context as a single entity, overlooking the fact that not all information provided in the context is important. Meanwhile, another line of work highlights the crucial role of attention mechanisms in identifying key contextual information, leveraging attention patterns for KV cache compression Li et al. (2024a) and document reordering Peysakhovich and Lerer (2023). Motivated by these insights, we delve deeper into the capabilities of attention-based mechanisms for fine-grained, inference-time evidence highlighting, enhancing LMs’ ability to concentrate on key supporting information within the context. Our contributions in this work are as follows:
@@
  • Leveraging this inherent ability, we propose an inference time context augmentation approach, SelfElicit, that highlights the important evidence in the context and improves LM’s ability to provide correct and well-grounded responses. SelfElicit is efficient for inference, training-free, and robust to noise and hyper-parameter choice.
  • In a comprehensive study, we show the significant performance improvement that SelfElicit brings across model families and tasks. We also offer extensive studies and analyses on various characteristics of SelfElicit, including but not limited to its evidence-eliciting accuracy, effect of different design choices, robustness under noisy/informative contexts, runtime efficiency, etc.
Fig. 1 illustrates the SelfElicit workflow. By identifying and highlighting key evidence within the context, SelfElicit guides the LM towards important information that it might otherwise overlook, thus leading to the correct response.
@@ 3.2 Contextual Evidence Highlighting
The overall procedure is summarized in Algorithm 1. We note that SelfElicit is designed to prioritize efficiency and simplicity, generating only one additional token during inference to identify ${\mathcal{S}}_{\text{SE}}$ (line 1 in Alg. 1). The highlighting method also outperforms more complex strategies, detailed further in Appendices B.5 and B.7.
@@ 4 Experiments
  • RQ1: How does SelfElicit perform in terms of improving answer quality and accuracy?
@@
Llama-3.1 8B Base 58.9 57.7 64.3 56.2 72.8 66.1 59.7 61.6 4.38 4.75 224.1
Cot 60.4 58.6 64.9 55.8 74.4 67.4 59.6 62.1 3.75 4.00 224.8
FullElicit 60.7 59.0 65.9 56.6 72.8 66.3 61.1 62.5 3.12 3.25 226.3
PromptElicit 66.3 66.3 62.8 56.7 76.0 69.6 61.8 65.6 2.75 2.00 1672.0
@@
70B Base 71.8 74.2 66.7 57.4 78.0 71.2 59.3 63.2 3.25 3.00 1389.8
Cot 72.4 74.0 67.2 56.4 77.0 69.9 60.3 63.1 3.12 4.50 1394.2
FullElicit 71.3 73.8 66.2 56.8 77.5 70.7 58.2 61.8 4.25 4.50 1408.1
PromptElicit 73.4 76.2 63.1 58.2 77.0 72.3 64.0 68.5 3.38 2.00 8124.0
@@
PromptElicit 71.0 64.1 57.7 42.2 81.9 62.3 65.6 42.8 4.00 1.25 1877.8
SelfElicit 74.3 61.6 60.6 41.7 83.6 61.3 66.4 43.4 2.00 1.75 431.3
12B Base 59.2 65.9 51.9 51.6 72.9 68.7 54.8 61.6 5.00 4.75 281.9
@@
FullElicit 59.8 65.8 53.1 51.9 73.7 68.8 55.5 62.6 2.88 4.00 283.9
PromptElicit 62.8 73.1 52.2 56.6 79.9 77.6 55.5 65.7 2.38 1.75 1455.0
SelfElicit 63.6 72.9 54.9 58.6 82.6 79.9 55.3 66.0 1.50 1.25 339.1
@@
Cot 70.7 37.9 59.2 31.9 78.6 41.6 63.8 32.3 2.00 5.00 421.5
FullElicit 65.5 65.7 57.5 48.2 77.1 67.3 64.4 60.6 3.50 2.75 249.6
PromptElicit 64.7 67.7 54.9 46.3 75.4 66.8 64.5 65.0 4.25 2.25 1165.1
SelfElicit 69.1 71.4 59.6 50.8 78.1 67.8 65.0 64.7 1.50 1.25 289.4
32B Base 71.8 68.4 60.0 44.7 79.0 69.3 62.7 59.3 3.12 3.25 928.2
Cot 71.3 67.1 60.0 43.5 79.5 66.8 59.5 55.3 3.62 5.00 998.6
FullElicit 71.3 68.2 61.6 45.7 78.5 68.8 63.2 58.1 3.25 3.75 936.3
PromptElicit 71.3 74.5 59.0 51.5 78.0 69.9 64.8 68.1 4.00 2.00 5109.8
@@ 4.1 Experimental Setup
  • Cot Wei et al. (2022a): Chain-of-thought prompting encourages the model to reason through intermediate steps before reaching a final answer. While CoT promotes step-by-step reasoning based on the evidence in the context, it does not explicitly highlight important information within the context. Therefore, we use it as a natural baseline for validating the advantage of explicit evidence elicitation in SelfElicit. We implement it by adding the Cot prompt "Think step by step to provide the answer." at the end of the instruction.
  • FullElicit: A naive approach that highlights the entire context as important. Comparing with it shows the necessity of fine-grained, sentence-level evidence elicitation.
@@ 4.2 Main Results
  • Cot also does not provide a meaningful and consistent gain which highlights the fact that asking the model to reason carefully does not improve its ability to leverage the relevant evidence while generating the response. Cot results also show variation across LM families where we don’t observe meaningful gains for Llama and Mistral models. When combined with Qwen models, Cot generates significantly longer answers with intermediate steps. While this results in a slight increase in EM scores in a few cases, it also leads to a large drop (up to 27.9) in F1 scores due to the inclusion of redundant information, as well as longer inference times (e.g., 171.9% more inference time for Qwen2.5-7B).
@@ 4.3 Additional Analysis
However, similar to the baseline, the performance of SelfElicit for "distractor" variant worsens meaningfully when compared to SelfElicit performance for the gold context setup, also shown in Figure 3(a). To investigate this we do a deep dive into the evidence elicitation capability of SelfElicit in the presence of substantial noise. Figure 3(b) contrasts the elicit ratio (i.e., the proportion of elicited evidence within the original context) under gold and distractor settings. With gold context, SelfElicit tends to select a larger proportion as evidence since most contextual information provided is relevant. By contrast, in the distractor setting where most of the context is irrelevant, the proportion of evidence selected decreases substantially (typically <10%) for the same $\alpha=0.5$. This shows that even with a static $\alpha$, SelfElicit naturally exhibits adaptiveness by selecting a lower evidence ratio in noisy contexts, focusing only on the relevant evidence without distraction from noise.
RQ4: Deeper Layers are better for ${\mathcal{L}}_{\text{ER}}$.
@@ 5 Related Works
Furnishing LMs with relevant context is an effective way of providing up-to-date external and/or private knowledge Ji et al. (2023); Asai et al. (2023) to help mitigate hallucination and improve response accuracy Huang et al. (2023). Retrieval-augmented Generation (RAG) is a widely adopted paradigm for this purpose Gao et al. (2023); Fan et al. (2024). Despite its popularity, recent studies have pointed out that context retrieved from external sources often contains noise and irrelevant information, leading to confusion for LM Cuconasu et al. (2024); Wu et al. (2024); Yu et al. (2024). Motivated by this, we explore how to leverage contextual information more effectively at a finer granularity by highlighting critical information within the context. To the best of our knowledge, we are the first to investigate automated contextual evidence highlighting based on LM internal representations.
@@
Recent studies have explored ways to understand QA factuality by analyzing the internal representations to identify important attention layers Yuksekgonul et al. (2024); Chen et al. (2024b) or heads Halawi et al. (2024) that are crucial for generation correctness or hallucination. However, they primarily focus on how LMs utilize their parametric knowledge in controlled generation Yuksekgonul et al. (2024); Halawi et al. (2024) or closed-book QA Chen et al. (2024b), with little discussion on context-based QA that rely on using external non-parametric knowledge. For context-based QA, Peysakhovich and Lerer (2023) propose combating positional bias by iteratively ranking retrieved context documents based on average attention scores. Wang et al. (2023b) employs a fine-tuned Llama-2 evidence extractor to filter context information. Nevertheless, these works either employ document-level average attention over all layers and heads or necessitate training an auxiliary model for generative filtering. In contrast, our work delves into a detailed analysis of the behavior of each attention layer at the sentence level, enabling fine-grained and efficient evidence highlighting at inference time.
@@ 6 Conclusion
We evaluated the effectiveness of SelfElicit across several open-source LMs. However, we can not assess SelfElicit on proprietary LMs due to the need to access attention scores. Prioritizing efficiency and simplicity, we used a static eliciting threshold in the paper. We discussed the effect of different threshold $\alpha$ choices in Figure 4. Our method is generally robust to different threshold $\alpha$, with a broad range of values yielding consistently strong improvements, but one can still fine-tune it on specific tasks for better performance. Moreover, an interesting future direction would be to automatically select the optimal elicit thresholds based on the model and input characteristics. We provide a more detailed case analysis and discussion on the potential benefit of dynamic threshold adjustment in Appendix C.1.
Acknowledgments
We thank Danai Koutra for her insights and discussions throughout the project. This work is partially supported by NSF (2134079 ) and AFOSR (FA9550-24-1-0002), The content of the information in this document does not necessarily reflect the position or the policy of the Government, and no official endorsement should be inferred. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation here on.
@@ Appendix A Reproducibility Details
We use PyTorch Paszke et al. (2019) and HuggingFace Transformers Wolf et al. (2020) to implement all involved language models (LMs). Specifically, the LMs used in the paper are the instruct fine-tuned version of Llama3.1 8B [^3] https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct and 70B [^4] https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct, Mistral 7B (instruct-v0.3) [^5] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3 and 12B (Nemo) [^6] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407, as well as Qwen2.5 7B [^7] https://huggingface.co/Qwen/Qwen2.5-7B-Instruct and 32B [^8] https://huggingface.co/Qwen/Qwen2.5-32B-Instruct. We use the official apply_chat_template function to get the input tokens from text input ($\tau({\mathbf{c}},{\mathbf{q}})$) and let the model generate the answer based on it. We use greedy decoding with a temperature of 0 to get deterministic answers.
Dataset details.
For all the datasets used, we test the LM on the official public validation split, with 7,405/4,212/7,785/12,836 context-question pairs for HotopotQA/NewsQA/TQA/NQ. The context passages in those datasets are retrieved from diverse sources including Wikipedia Yang et al. (2018); Kwiatkowski et al. (2019), news articles Trischler et al. (2017), and web search queries Joshi et al. (2017). The detailed statistics and descriptions are reported in Table 5. We use Llama-3.1 tokenizer to compute the number of tokens in the context passage.
Table 5: Dataset description and statistics.
Dataset #ContextTokens #Samples Source
Avg. Max
HotpotQA 1251.6 3346 7405 Wikipedia
NewsQA 625.24 940 4212 CNN News Article
TQA 892.89 1974 7785 Bing Search Query
NQ 249.38 2679 12836 Wikipedia
Table 6: Detailed runtime and efficiency results. we report the average inference time and number of generated tokens for each sample (i.e., context-question pair) for each dataset.
Model Method Dataset Average
HotpotQA NewsQA TQA NQ Inference #Generated
Time (ms) #Tokens Time (ms) #Tokens Time (ms) #Tokens Time (ms) #Tokens Time/sample (ms) Tokens
Llama-3.1 8B Base 161.46 6.68 290.76 11.79 230.18 6.72 213.18 10.42 223.90 (+0.0%) 8.90 (+0.0%)
Cot 163.14 6.52 290.86 11.77 230.21 6.76 213.85 10.36 224.51 (+0.3%) 8.85 (-0.6%)
FullElicit 163.55 6.49 290.79 11.83 236.30 6.97 213.64 10.47 226.07 (+1.0%) 8.94 (+0.4%)
PromptElicit 1419.08 75.78 1718.02 85.18 2412.91 116.81 1117.64 61.40 1666.91 (+644.5%) 84.79 (+852.5%)
SelfElicit 204.65 6.98 318.74 10.54 297.71 7.44 233.78 9.83 263.72 (+17.8%) 8.70 (-2.3%)
70B Base 910.68 5.17 1850.04 11.32 1397.99 5.95 1421.32 11.18 1395.01 (+0.0%) 8.40 (+0.0%)
Cot 917.37 5.15 1856.47 10.61 1398.79 5.88 1423.61 10.88 1399.06 (+0.3%) 8.13 (-3.2%)
FullElicit 915.09 5.41 1851.00 11.13 1448.56 6.32 1434.97 10.31 1412.41 (+1.2%) 8.29 (-1.3%)
PromptElicit 8545.64 69.39 7712.69 55.77 8283.76 56.14 7139.59 60.06 7920.42 (+467.8%) 60.34 (+618.1%)
SelfElicit 1198.53 6.05 1854.49 9.48 1672.19 6.15 1398.81 9.88 1531.00 (+9.7%) 7.89 (-6.1%)
Mistral 7B Base 468.26 21.60 564.08 24.60 479.97 18.67 647.02 31.38 539.83 (+0.0%) 24.06 (+0.0%)
Cot 489.55 23.31 568.91 25.27 502.71 19.59 688.56 34.02 562.43 (+4.2%) 25.55 (+6.2%)
FullElicit 477.01 19.74 567.08 23.57 479.99 18.61 647.23 31.78 542.83 (+0.6%) 23.42 (-2.7%)
PromptElicit 1444.44 70.58 2043.11 97.09 2230.07 102.03 1788.88 89.59 1876.62 (+247.6%) 89.82 (+273.3%)
SelfElicit 330.42 12.74 493.56 18.33 440.59 13.32 462.43 21.00 431.75 (-20.0%) 16.34 (-32.1%)
12B Base 205.88 5.62 380.22 10.46 287.35 5.32 251.74 7.97 281.30 (+0.0%) 7.34 (+0.0%)
Cot 208.52 5.57 381.90 10.09 287.36 5.35 258.02 7.41 283.95 (+0.9%) 7.11 (-3.2%)
FullElicit 208.39 5.65 380.59 10.54 289.87 5.41 254.23 7.90 283.27 (+0.7%) 7.37 (+0.4%)
PromptElicit 1273.57 46.77 1560.04 52.11 1749.14 54.64 1226.14 46.18 1452.22 (+416.3%) 49.93 (+580.0%)
SelfElicit 278.07 6.29 428.54 9.02 385.85 5.45 260.08 7.47 338.14 (+20.2%) 7.06 (-3.9%)
Qwen2.5 7B Base 143.78 6.36 303.72 13.53 247.26 8.26 286.31 15.49 245.27 (+0.0%) 10.91 (+0.0%)
Cot 449.74 25.10 407.30 20.06 370.10 15.93 460.24 26.51 421.85 (+72.0%) 21.90 (+100.7%)
FullElicit 145.24 6.18 306.38 12.89 255.94 6.85 291.09 16.35 249.66 (+1.8%) 10.57 (-3.1%)
PromptElicit 1109.49 62.77 1227.45 64.27 1312.63 64.84 1005.11 58.32 1163.67 (+374.4%) 62.55 (+473.3%)
SelfElicit 194.21 6.90 357.74 12.92 292.41 7.59 313.18 16.05 289.38 (+18.0%) 10.87 (-0.4%)
32B Base 624.64 6.55 1271.39 15.35 773.30 6.51 1044.09 14.50 928.35 (+0.0%) 10.73 (+0.0%)
Cot 625.01 7.46 1447.84 18.24 871.12 8.13 1048.22 14.59 998.05 (+7.5%) 12.10 (+12.8%)
FullElicit 636.89 6.63 1275.19 15.06 780.95 6.80 1052.68 14.21 936.43 (+0.9%) 10.67 (-0.5%)
PromptElicit 5556.76 82.86 5468.44 76.23 5290.17 69.83 4096.52 62.62 5102.97 (+449.7%) 72.89 (+579.5%)
SelfElicit 738.98 6.32 1220.44 12.29 1011.58 7.02 948.09 11.47 979.77 (+5.5%) 9.28 (-13.5%)